# The Invisible Hand of AI Libraries Shaping Open Source Projects and Communities

### v1

@@ I Introduction
In the early 1980s, Open Source Software (OSS) emerged as a revolutionary concept amidst the dominance of proprietary software. To keep the OSS momentum, Richard Stallman, in 1985, authored the GNU Manifesto [stallman1985gnu], marking a pivotal moment in the OSS movement. What began as a revolutionary idea [doi:10.1080/08874417.2009.11646043] has now become the cornerstone of computer science [1541831], with OSS finding widespread adoption across both academic research and industry [8737777]. Artificial intelligence (AI) thrives within the realm of OSS. The widespread of libraries capable of managing sophisticated Machine Learning (ML) algorithms or designing intricate neural networks has accelerated advancements in both research and industry, paving the way to “software 2.0[10.1145/3453478].
State-of-the-art commercial and OSS products expand their features daily by integrating AI libraries into OSS [DBLP:journals/tse/YuBSM21]. AI libraries, comprising pre-built algorithms and frameworks, enable developers to implement AI functionalities in their projects seamlessly [10274712]. AI libraries have democratized access to AI capabilities, enabling developers of varying skill levels to incorporate AI-driven features into their software projects without extensive machine learning or AI development expertise [9695219]. Hence, it fosters inclusivity within the developer community, allowing individuals from diverse backgrounds to contribute to and benefit from AI-driven innovation [10274712, 9695219, DBLP:journals/tse/YuBSM21]. Developers leverage AI libraries as building blocks, accelerating the development cycle, improving governance, and fostering a culture of innovation within open-source projects [9695219, esposito2025icseseip].
@@ II Background & Related Work
One distinguishing feature of OSS is that its source code is openly accessible for anyone to use, modify, and distribute. In contrast to proprietary software developed and owned by individual companies, OSS fosters collaborative and community-driven development Fortunato et al. [fortunato2021case]. These characteristics make OSS a suitable candidate for diverse research in software engineering Gyimothy et al. [1542070].
@@ II-A OSS in Empirical Software Engineering
Mockus et al. [10.1145/567793.567795] compare commercial and open-source methodologies using Mozilla and Apache as examples. They find that while commercial projects rely on explicit coordination, OSS projects like Apache use decentralized communication. Mozilla faces challenges due to module interdependencies, while Apache’s approach emphasizes modularity. Leveraging the open-source community for low-interdependence tasks can improve productivity.
Similarly, Gyimothy et al. [1542070] conducted a study to assess the quality and reliability of OSS systems, mainly focusing on Mozilla, a widely used web and email suite. They calculated object-oriented metrics proposed by Chidamber et al. [chidamber1994metrics] to analyze fault-proneness in Mozilla’s source code. Moreover, Yu et al. [DBLP:journals/tse/YuBSM21] conducted a large-scale empirical study on Java annotations, analyzing 1,094 OSS projects on GitHub. They investigated annotation usage, evolution, and impact, revealing 10 novel findings. Similarly, we are also interested in the context of OSS projects. Nonetheless, the author focused on Java annotations, while our study focuses on AI libraries and technologies.
@@
Esposito et al. [esposito2023uncovering] investigated the hidden bugs, i.e., buggy untouched methods, that remained dormant for various releases of OSS projects. Similarly, Esposito et al. [esposito2023can] performed a large-scale analysis on OSS vulnerability severity according to SonarQube and the National Vulnerability (NVD) Datasets to assess whether the NVD vulnerability severity was correlated with SQ severity levels. Finally, Esposito et al. [esposito2024extensive] performed a large-scale OSS-oriented test of static analysis security testing tools. Moreover, recent studies employed the ready availability of OSS development history to improve JIT defect prediction
@@ II-B AI libraries in Software Engineering
Moran et al. [8374985] devised an automated approach called ReDraw to streamline the transformation of GUI mock-ups into code for user-facing software. ReDraw utilizes computer vision techniques and deep convolutional neural networks. A data-driven algorithm then generates a hierarchical GUI structure, enabling automatic assembly of prototype applications for the Android platform. Esposito et al. [esposito2024validate], compiled a systematic dataset review on the usage of AI in the field of vulnerability prediction.
@@
Xu et al. [10.1145/3487569], address the challenge of turning conceptual ideas into code, especially when dealing with unfamiliar library APIs, by investigating the promise and challenges of using ML for code generation and retrieval within the PyCharm IDE. They developed a plugin for PyCharm that combines code generation and retrieval functionality and orchestrates virtual environments to collect user events. While qualitative surveys indicate positive developer experiences, quantitative results regarding increased productivity, code quality, and program correctness are inconclusive. The study identifies critical points for improving future machine learning-based code assistants and demonstrates developers’ preferences between code generation and retrieval.
@@ II-C AI libraries in OSS
Dilhara et al. [10.1145/3453478] presented a large-scale empirical study to examine the challenges developers face utilizing ML libraries in Software-2.0 development. Key findings include a significant increase in the adoption of ML libraries, various usage patterns, and unique challenges such as the binary incompatibility of trained ML models. Our study focuses on the broader AI spectrum of libraries, evaluating the impact of adopting them.
Tufano et al. [tufano2024unveiling] investigated how developers use Large Language Models, like OpenAI’s ChatGPT, in OSS projects. Out of the 1,501 instances of potential ChatGPT usage in open-source projects, the authors narrowed them down to 467 actual ChatGPT instances. [tufano2024unveiling] categorized these instances into 45 tasks developers automate using ChatGPT, hence helping developers understand how to use LLMs effectively and providing researchers with insight into tasks that can benefit and those that do not from automated solutions. Our study focuses on the usage of AI libraries, thus including LLMs, in OSS projects, intending to investigate the impact that the adoption of such libraries has on community engagement and the technical ecosystem.
@@ III Empirical Study Design
Our empirical study design follows the guidelines defined by Wohlin et al. [Wohlin2000]. In the upcoming sections, we detail the specific research questions and goals that drive our study and the procedures we used for data collection and analysis.
@@ III-A Goal, Research Questions, and Hypothesis
We formalized the goal of this study according to the Goal Question Metric (GQM) approach [Basili1994] as follows:
@@
This $RQ$ aims to measure the extent of AI library adoption in OSS projects. AI adoption captures whether and how much a project integrates AI features and capabilities through external components, i.e., libraries. According to [tarrega2020measuring, kula2018developers, pashchenko2018vulnerable], we can analyze evidence of a project dependency on three levels:
@@
In this first $RQ$, we aim to measure the dependency-level evidence, specifically whether the OSS project includes references to AI libraries in its list of dependencies. Furthermore, we are keen to focus on Java and Python because they represent two dominant yet contrasting ecosystems in software development [cass2025top-programming-languages]. Python is the de facto language for AI and machine learning, offering a rich ecosystem of libraries, including TensorFlow, PyTorch, and scikit-learn [li2025bridging, khandare2023analysis]. In contrast, Java is widely adopted in enterprise and large-scale software systems, where AI integration is emerging [Azul2025]. Studying both allows us to capture differences between experimentation-oriented (Python) and production-oriented (Java) environments. Other languages were excluded as they either lack mature AI ecosystems (e.g, Go) [JetBrainsGoML2023, JetBrainsGoTrends2024], or they are more challenging, e.g., steeper learning curve, memory management complexity, that may affect adoption compared to higher-level languages (e.g. C++) [GeeksforGeeksCPlusPlusML], or are less representative of large, diverse open-source communities suitable for comparative analysis [li2025bridging].
@@
In recent years, OSS projects have increasingly embraced AI libraries like TensorFlow, PyTorch, and Hugging Face to build smarter, data-driven applications covering a wide range of possible application fields spanning from innovative products to code generation and co-architecting complex software solutions [chiu2023impact, makridakis2017forthcoming, dwivedi2021artificial, esposito_generative_2025]. Albeit such libraries bring powerful capabilities, they may also change how teams work, communities engage, and technology choices are made [akbar_6gsoft_2024]. While AI libraries are becoming a staple in modern development, we still know surprisingly little about how they shape the projects that adopt them. Do they influence how developers collaborate? Do they affect how often code is released or how issues and pull requests are handled? Understanding whether adopting AI libraries may shape some project aspects is crucial, especially as more OSS communities move toward AI-enabled development [lagrandeur2023consequences]. Therefore, we aim to explore how adopting AI libraries impacts OSS projects across three key dimensions: how teams collaborate, how the community engages, and how technologies and workflows evolve. Following previous studies, we collected repository metrics (RMs) that provide valuable insights into the activity, health, and community engagement of a software project hosted on platforms like GitHub to measure such key aspects. Due to space constraints, we present in Table I (online appendix) the collectable RMs, their descriptions, and their classifications according to [d2010extensive, bavota2015four, bugayenko2024cam, github_docs]. Based on our RQ goal and the RM metrics, we conjecture the following seven null-hypotheses ($H_{0}$) and alternative hypotheses ($H_{1}$):
@@
AI libraries often bring new architectural choices, advanced abstractions, and dense logic, which may shape development practices in subtle but impactful ways [esposito2024leveraging]. While much attention has been paid to the capabilities these libraries enable, we still know little about their side effects on internal software quality. Do they make code more complex? Less cohesive? Do they change the way developers design and organize classes?
With open-source repositories playing a central role in shaping modern AI-enabled applications, these questions are more relevant than ever. With this study, we aim to explore whether, and how, the inclusion of AI libraries influences structural, complexity-related, and maintainability aspects of source code in open-source projects. To do this, we group our analysis around three core dimensions: (i) object-oriented design structure, (ii) control-flow complexity and cohesion, and (iii) maintainability and readability of the codebase. The goal is to identify statistically significant patterns that distinguish AI-enhanced projects from traditional ones at a deeper, code-centric level. Hence, to measure changes in such dimensions, we opted to employ software product and quality metrics (SMs) to provide quantitative insights into various aspects of software design, implementation, and maintainability [esposito2023uncovering], obtainable via Understand [scitools_understand_2025] (Table II in the Online Appendix). Based on our RQ goal and the SMs, we conjecture the following three null-hypotheses ($H_{0}$) and alternative hypotheses ($H_{1}$):
@@ IV-A Study Setup and Data Collection
This section describes our data collection methodology. Figure 1 depicts our data collection workflow that utilizes the BPMN 2.0 notation [object_management_group_business_2011]. According to Figure 1, our approach starts by programmatically accessing GitHub repositories to gather the raw data required for our study. We will employ GitHub Search and Repository APIs, performing two independent searches requesting repositories with Java or Python as their primary programming language.
@@ IV-A1 GitHub Project Selection
The GitHub repository API offers a seamless way to interact with repositories without downloading them. Similar to [dabic2021sampling], we requested all repositories created in a specific time to overcome the API limitation. We define a dynamic time window. We will analyze the repositories created between the 1st of January 2000, and the data collection date. We default to using a time window of five years. In the case of precisely 1000 results, we reduced the window to one year, six months, one month, and one week [dabic2021sampling]. [10.1145/3453478, DBLP:journals/tse/YuBSM21] inspired our selection process. Table I summarizes the inclusion criteria for repositories. According to Table I, we will retrieve all repositories whose star count (SC) will be greater than 50 when performing the full study [10.1145/3453478, DBLP:journals/tse/YuBSM21]. We employ SC as a proxy metric for repository popularity according to [BORGES2018112]. We will include only non-forked and unarchived projects and avoid inactive ones by discarding those that do not have a single activity event within six months before our data collection date [10.1145/3453478]. Moreover, we included only projects that allowed access to the Software Bill of Materials [^1] https://docs.github.com/en/code-security/supply-chain-security/understanding-your-software-supply-chain/about-the-dependency-graph (SBOM) to obtain the dependency list. Finally, according to [BORGES2018112], we exclude repositories with fewer than three contributors.
@@ IV-A2 AI Libraries Adoption ($RQ_{1}$)
Leveraging the final project selection, we collect the dependencies list to measure the adoption of AI libraries in OSS projects ($RQ_{1}$). More specifically, we will leverage the SBOMs, a structured inventory that provides comprehensive details about a software project’s components, to analyze the adoption of AI libraries. It includes component names, versions, dependencies, licensing details, and sometimes security vulnerabilities [10.1145/3453478]. Parsing the SBOM files allows us to gather the list of all dependencies of each repository, enabling our analysis.
@@ IV-A3 Selection of AI Libraries
For each candidate library, we collect popularity metrics as a proxy of widespread adoption [bavota2021dependencies, dilhara2021understanding]: for PyPI, we collect download counts; for Maven Central, when available, download and reverse-dependency counts. In the case that Maven libraries do not have such information, we collect from the library’s GitHub repository the stars count metric. We compute the interquartile range (IQR) for the primary popularity metrics (stars for GitHub, downloads for registries) and retain libraries at or above $Q_{3}$ therefore removing long-tail entries with limited adoption, yielding a focused set of potentially widely used AI libraries for analysis.
Finally, we manually validate the shortlisted libraries to confirm their genuine AI orientation. Two authors will independently examine documentation, README files, and dependency structures to verify that each library actually provides AI features, thereby excluding “AI-assisting” libraries such as NumPy or Pandas that may be associated with AI workflows but do not provide any actual AI features. We measure the agreement via Cohen’s $\kappa$ [cohen_coefficient_1960] and solve disagreements involving the third author. Finally, for each AI library, we collect its package identifiers and match against the SBOM for each project.

### v2

@@ I Introduction
In the early 1980s, Open Source Software (OSS) emerged as a revolutionary concept amidst the dominance of proprietary software. To keep the OSS momentum, Richard Stallman, in 1985, authored the GNU Manifesto [61], marking a pivotal moment in the OSS movement. What began as a revolutionary idea [12] has now become the cornerstone of computer science [38], with OSS finding widespread adoption across both academic research and industry [11]. Artificial intelligence (AI) thrives within the realm of OSS. The widespread of libraries capable of managing sophisticated Machine Learning (ML) algorithms or designing intricate neural networks has accelerated advancements in both research and industry, paving the way to “software 2.0[21].
State-of-the-art commercial and OSS products expand their features daily by integrating AI libraries into OSS [72]. AI libraries, comprising pre-built algorithms and frameworks, enable developers to implement AI functionalities in their projects seamlessly [1]. AI libraries have democratized access to AI capabilities, enabling developers of varying skill levels to incorporate AI-driven features into their software projects without extensive machine learning or AI development expertise [3]. Hence, it fosters inclusivity within the developer community, allowing individuals from diverse backgrounds to contribute to and benefit from AI-driven innovation [1, 3, 72]. Developers leverage AI libraries as building blocks, accelerating the development cycle, improving governance, and fostering a culture of innovation within open-source projects [3, 30].
@@ II Background & Related Work
One distinguishing feature of OSS is that its source code is openly accessible for anyone to use, modify, and distribute. In contrast to proprietary software developed and owned by individual companies, OSS fosters collaborative and community-driven development Fortunato et al. [35]. These characteristics make OSS a suitable candidate for diverse research in software engineering Gyimothy et al. [39].
@@ II-A OSS in Empirical Software Engineering
Mockus et al. [50] compare commercial and open-source methodologies using Mozilla and Apache as examples. They find that while commercial projects rely on explicit coordination, OSS projects like Apache use decentralized communication. Mozilla faces challenges due to module interdependencies, while Apache’s approach emphasizes modularity. Leveraging the open-source community for low-interdependence tasks can improve productivity.
Similarly, Gyimothy et al. [39] conducted a study to assess the quality and reliability of OSS systems, mainly focusing on Mozilla, a widely used web and email suite. They calculated object-oriented metrics proposed by Chidamber et al. [15] to analyze fault-proneness in Mozilla’s source code. Moreover, Yu et al. [72] conducted a large-scale empirical study on Java annotations, analyzing 1,094 OSS projects on GitHub. They investigated annotation usage, evolution, and impact, revealing 10 novel findings. Similarly, we are also interested in the context of OSS projects. Nonetheless, the author focused on Java annotations, while our study focuses on AI libraries and technologies.
@@
Esposito et al. [26] investigated the hidden bugs, i.e., buggy untouched methods, that remained dormant for various releases of OSS projects. Similarly, Esposito et al. [29] performed a large-scale analysis on OSS vulnerability severity according to SonarQube and the National Vulnerability (NVD) Datasets to assess whether the NVD vulnerability severity was correlated with SQ severity levels. Finally, Esposito et al. [25] performed a large-scale OSS-oriented test of static analysis security testing tools. Moreover, recent studies employed the ready availability of OSS development history to improve JIT defect prediction
@@ II-B AI libraries in Software Engineering
Moran et al. [52] devised an automated approach called ReDraw to streamline the transformation of GUI mock-ups into code for user-facing software. ReDraw utilizes computer vision techniques and deep convolutional neural networks. A data-driven algorithm then generates a hierarchical GUI structure, enabling automatic assembly of prototype applications for the Android platform. Esposito et al. [27], compiled a systematic dataset review on the usage of AI in the field of vulnerability prediction.
@@
Xu et al. [71], address the challenge of turning conceptual ideas into code, especially when dealing with unfamiliar library APIs, by investigating the promise and challenges of using ML for code generation and retrieval within the PyCharm IDE. They developed a plugin for PyCharm that combines code generation and retrieval functionality and orchestrates virtual environments to collect user events. While qualitative surveys indicate positive developer experiences, quantitative results regarding increased productivity, code quality, and program correctness are inconclusive. The study identifies critical points for improving future machine learning-based code assistants and demonstrates developers’ preferences between code generation and retrieval.
@@ II-C AI libraries in OSS
Dilhara et al. [21] presented a large-scale empirical study to examine the challenges developers face utilizing ML libraries in Software-2.0 development. Key findings include a significant increase in the adoption of ML libraries, various usage patterns, and unique challenges such as the binary incompatibility of trained ML models. Our study focuses on the broader AI spectrum of libraries, evaluating the impact of adopting them.
Tufano et al. [65] investigated how developers use Large Language Models, like OpenAI’s ChatGPT, in OSS projects. Out of the 1,501 instances of potential ChatGPT usage in open-source projects, the authors narrowed them down to 467 actual ChatGPT instances. [65] categorized these instances into 45 tasks developers automate using ChatGPT, hence helping developers understand how to use LLMs effectively and providing researchers with insight into tasks that can benefit and those that do not from automated solutions. Our study focuses on the usage of AI libraries, thus including LLMs, in OSS projects, intending to investigate the impact that the adoption of such libraries has on community engagement and the technical ecosystem.
@@ III Empirical Study Design
Our empirical study design follows the guidelines defined by Wohlin et al. [69]. In the upcoming sections, we detail the specific research questions and goals that drive our study and the procedures we used for data collection and analysis.
@@ III-A Goal, Research Questions, and Hypothesis
We formalized the goal of this study according to the Goal Question Metric (GQM) approach [7] as follows:
@@
This $RQ$ aims to measure the extent of AI library adoption in OSS projects. AI adoption captures whether and how much a project integrates AI features and capabilities through external components, i.e., libraries. According to [64, 44, 55], we can analyze evidence of a project dependency on three levels:
@@
In this first $RQ$, we aim to measure the dependency-level evidence, specifically whether the OSS project includes references to AI libraries in its list of dependencies. Furthermore, we are keen to focus on Java and Python because they represent two dominant yet contrasting ecosystems in software development [14]. Python is the de facto language for AI and machine learning, offering a rich ecosystem of libraries, including TensorFlow, PyTorch, and scikit-learn [46, 42]. In contrast, Java is widely adopted in enterprise and large-scale software systems, where AI integration is emerging [6]. Studying both allows us to capture differences between experimentation-oriented (Python) and production-oriented (Java) environments. Other languages were excluded as they either lack mature AI ecosystems (e.g, Go) [40, 41], or they are more challenging, e.g., steeper learning curve, memory management complexity, that may affect adoption compared to higher-level languages (e.g. C++) [36], or are less representative of large, diverse open-source communities suitable for comparative analysis [46].
@@
In recent years, OSS projects have increasingly embraced AI libraries like TensorFlow, PyTorch, and Hugging Face to build smarter, data-driven applications covering a wide range of possible application fields spanning from innovative products to code generation and co-architecting complex software solutions [16, 48, 24, 28]. Albeit such libraries bring powerful capabilities, they may also change how teams work, communities engage, and technology choices are made [4]. While AI libraries are becoming a staple in modern development, we still know surprisingly little about how they shape the projects that adopt them. Do they influence how developers collaborate? Do they affect how often code is released or how issues and pull requests are handled? Understanding whether adopting AI libraries may shape some project aspects is crucial, especially as more OSS communities move toward AI-enabled development [45]. Therefore, we aim to explore how adopting AI libraries impacts OSS projects across three key dimensions: how teams collaborate, how the community engages, and how technologies and workflows evolve. Following previous studies, we collected repository metrics (RMs) that provide valuable insights into the activity, health, and community engagement of a software project hosted on platforms like GitHub to measure such key aspects. Due to space constraints, we present in Table I (online appendix) the collectable RMs, their descriptions, and their classifications according to [18, 8, 10, 37]. Based on our RQ goal and the RM metrics, we conjecture the following seven null-hypotheses ($H_{0}$) and alternative hypotheses ($H_{1}$):
@@
AI libraries often bring new architectural choices, advanced abstractions, and dense logic, which may shape development practices in subtle but impactful ways [31]. While much attention has been paid to the capabilities these libraries enable, we still know little about their side effects on internal software quality. Do they make code more complex? Less cohesive? Do they change the way developers design and organize classes?
With open-source repositories playing a central role in shaping modern AI-enabled applications, these questions are more relevant than ever. With this study, we aim to explore whether, and how, the inclusion of AI libraries influences structural, complexity-related, and maintainability aspects of source code in open-source projects. To do this, we group our analysis around three core dimensions: (i) object-oriented design structure, (ii) control-flow complexity and cohesion, and (iii) maintainability and readability of the codebase. The goal is to identify statistically significant patterns that distinguish AI-enhanced projects from traditional ones at a deeper, code-centric level. Hence, to measure changes in such dimensions, we opted to employ software product and quality metrics (SMs) to provide quantitative insights into various aspects of software design, implementation, and maintainability [26], obtainable via Understand [58] (Table II in the Online Appendix). Based on our RQ goal and the SMs, we conjecture the following three null-hypotheses ($H_{0}$) and alternative hypotheses ($H_{1}$):
@@ IV-A Study Setup and Data Collection
This section describes our data collection methodology. Figure 1 depicts our data collection workflow that utilizes the BPMN 2.0 notation [54]. According to Figure 1, our approach starts by programmatically accessing GitHub repositories to gather the raw data required for our study. We will employ GitHub Search and Repository APIs, performing two independent searches requesting repositories with Java or Python as their primary programming language.
@@ IV-A1 GitHub Project Selection
The GitHub repository API offers a seamless way to interact with repositories without downloading them. Similar to [19], we requested all repositories created in a specific time to overcome the API limitation. We define a dynamic time window. We will analyze the repositories created between the 1st of January 2000, and the data collection date. We default to using a time window of five years. In the case of precisely 1000 results, we reduced the window to one year, six months, one month, and one week [19]. [21, 72] inspired our selection process. Table I summarizes the inclusion criteria for repositories. According to Table I, we will retrieve all repositories whose star count (SC) will be greater than 50 when performing the full study [21, 72]. We employ SC as a proxy metric for repository popularity according to [9]. We will include only non-forked and unarchived projects and avoid inactive ones by discarding those that do not have a single activity event within six months before our data collection date [21]. Moreover, we included only projects that allowed access to the Software Bill of Materials [^1] https://docs.github.com/en/code-security/supply-chain-security/understanding-your-software-supply-chain/about-the-dependency-graph (SBOM) to obtain the dependency list. Finally, according to [9], we exclude repositories with fewer than three contributors.
@@ IV-A2 AI Libraries Adoption ($RQ_{1}$)
Leveraging the final project selection, we collect the dependencies list to measure the adoption of AI libraries in OSS projects ($RQ_{1}$). More specifically, we will leverage the SBOMs, a structured inventory that provides comprehensive details about a software project’s components, to analyze the adoption of AI libraries. It includes component names, versions, dependencies, licensing details, and sometimes security vulnerabilities [21]. Parsing the SBOM files allows us to gather the list of all dependencies of each repository, enabling our analysis.
@@ IV-A3 Selection of AI Libraries
For each candidate library, we collect popularity metrics as a proxy of widespread adoption [20, 22]: for PyPI, we collect download counts; for Maven Central, when available, download and reverse-dependency counts. In the case that Maven libraries do not have such information, we collect from the library’s GitHub repository the stars count metric. We compute the interquartile range (IQR) for the primary popularity metrics (stars for GitHub, downloads for registries) and retain libraries at or above $Q_{3}$ therefore removing long-tail entries with limited adoption, yielding a focused set of potentially widely used AI libraries for analysis.
Finally, we manually validate the shortlisted libraries to confirm their genuine AI orientation. Two authors will independently examine documentation, README files, and dependency structures to verify that each library actually provides AI features, thereby excluding “AI-assisting” libraries such as NumPy or Pandas that may be associated with AI workflows but do not provide any actual AI features. We measure the agreement via Cohen’s $\kappa$ [17] and solve disagreements involving the third author. Finally, for each AI library, we collect its package identifiers and match against the SBOM for each project.